# What Matters in Learning Facts in Language Models? Multifaceted Knowledge Probing with Diverse Multi-Prompt Datasets

### v1

@@ Abstract
Large language models (LLMs) face issues in handling factual knowledge, making it vital to evaluate their true ability to understand facts. In this study, we introduce knowledge probing frameworks, BELIEF(-ICL), to evaluate the knowledge understanding ability of both encoder-based and decoder-based pre-trained LMs (PLMs) from diverse perspectives. BELIEFs utilize a multi-prompt dataset to evaluate PLM’s accuracy, consistency, and reliability in factual knowledge understanding. To enable a more reliable evaluation with BELIEFs, we semi-automatically create MyriadLAMA, which has massively diverse prompts. We validate the effectiveness of BELIEFs in correctly and comprehensively evaluating PLM’s factual understanding ability via extensive evaluations with recent LLMs. We then investigate key factors in learning facts in LLMs, and reveal the limitation of the prompt-based knowledge probing. The dataset is anonymously publicized. [^1] https://anonymous.4open.science/r/belief-CC8A
What Matters in Learning Facts in Language Models?
Multifaceted Knowledge Probing with Diverse Multi-Prompt Datasets
Xin Zhao The University of Tokyo xzhao@tkl.iis.u-tokyo.ac.jp                        Naoki Yoshinaga   Daisuke Oba Institute of Industrial Science, The University of Tokyo {ynaga,oba}@iis.u-tokyo.ac.jp
@@ 1 Introduction
One of the strongest motivations for training a language model (LM) using a larger amount of text is to increase the LM’s ability to manage factual knowledge in knowledge-intensive tasks Kamalloo et al. (2023). However, even if LMs are trained on massive text, they suffer from hallucinations that generate incorrect knowledge-grounded sentences Zhang et al. (2023). Considering that large language models (LLMs) are being widely applied to real-world tasks, it is vital to understand the true knowledge capacity of LLMs and what factors influence the knowledge learning during pre-training.
@@
However, evaluating the LLM’s knowledge capacity is still challenging. Although LAMA probing Petroni et al. (2019) evaluates the knowledge contained in pre-trained LMs (PLMs), it provides only prediction accuracy. Some studies diversify the LAMA probing datasets to compute prediction consistency (robustness) Elazar et al. (2021); Jiang et al. (2020), but those datasets have either low quality or low quantity issues (§ B). Moreover, since the LAMA probe assumes encoder-based PLMs with the masked LM objective to solve fill-in-the-blank tasks, directly applying it to recent decoder-based LLMs will underestimate their ability of knowledge understanding. Although recent studies leveraged QA-based datasets to probe LLMs’ knowledge Kalo and Fichtel (2022); Mallen et al. (2023); Wiland et al. (2024); Maekawa et al. (2024) via prediction accuracy, they overlook other important aspects than accuracy such as robustness to diverse prompts and the reliability of predictions, which are important for real-world applications.
@@
In this study, we introduce BELIEF (§2) and its variant BELIEF-ICL (§3), the multifaceted factual knowledge probing frameworks for encoder- and decoder-based PLMs. BELIEFs utilize diverse prompts for each fact to eliminate the specific influence of linguistic expression on the results. It enables us to evaluate PLMs’ robustness and reliability by measuring accuracy fluctuations, consistency, and overconfidence in fact prediction. Since BELIEFs require a multi-prompt probing dataset with diverse prompts for each fact, we build a new probing dataset, MyriadLAMA, to enable a more accurate and comprehensive evaluation (§4).
@@
We applied BELIEFs to various encoder- and decoder-based PLMs, including BERT Devlin et al. (2019) and Llama2 Touvron et al. (2023)5.1). Through extensive evaluations, we verify the effectiveness of BELIEFs in accurately uncovering PLMs’ ability in factual understanding (§5). Moreover, by comparing different PLMs, we reveal insights into the factors affecting the acquisition of facts during pre-training (§6). Finally, we explore the upper limits of covered knowledge by different methods and reveal the limitations of prompt-based knowledge probing (§7).
2 BELIEF Framework
@@
We first present a multifaceted factual probing framework - BELIEF, specifically for encoder-based PLMs. Using a multi-prompt knowledge probing dataset, BELIEF evaluates the facts understanding ability of PLMs from accuracy, robustness and reliability (§2.2-2.4). Here, robustness measures PLMs’ ability to maintain consistent accuracy and predictions when given different prompts in evaluation. Reliability reflects the extent to which we can trust the PLMs’ predictions.
2.1 Preliminary
@@
To accurately evaluate the facts in PLMs, BELIEF aggregates results from multiple prompts for each fact to mitigate biases from specific linguistic expressions. This requires a dataset with varied expressions for each fact, namely multi-prompt factual probing dataset.
@@ 2.2 Accuracy and its fluctuations
Fluctuation of accuracy: For $V_{\textrm{Acc}@1}$, we evaluate accuracy fluctuations using the range and standard deviation. The range is determined by the difference between the maximum and minimum accuracy values in $V_{\textrm{Acc}@1}$.
@@ 2.4 Reliability
Overconf $=\sum_{i=1}^{M}\frac{|P^{(i)}|}{M}(\overline{o_{max}}^{(i)}-\overline{\textrm{Acc}@1}^{(i)})$ (2)
The closer the Overconf is to zero, the more aligned the model’s confidence is with its accuracy, indicating reliable confidence. A negative Overconf value means the model is underconfident.
@@ 3.1 In-context learning for fact probe
2) Context settings: We propose four types of contexts to assess the impact of examplar selection on factual knowledge probing, following the QA format outlined in InstructGPT Ouyang et al. (2022).
zero-shot: no context;
X-random: sampling X facts from all relations as the few-shot learning examples;
X-relation: sampling X facts from the same relation but with random templates;
X-template: sampling X facts from the same relations and the same template.
@@ 3.2 Evaluation methods
To measure BELIEF metrics for LLMs, we need to compare two strings: the generated text and the correct target expression for Acc@1, and two generated texts for Consist and Overconf. Here, we first normalize the strings by tokenizing and lemmatizing them. For example, "a guitar" and "guitars" are normalized to “a, guitar” and “guitar.” If a normalized string list is included in the other (partial matching), they are considered matched.
@@
3) Reliability: To calculate overconfidence, we need the model’s confidence (probability) in its output. However, we cannot obtain this directly from the probability over generated tokens, as LLMs can produce diverse outputs that represent the same answer. To address this, we propose an approximate measurement. For each prompt, we generate 100 samples using multinomial sampling [^2] Multinomial sampling selects a next token according to the probability, over the entire vocabulary given by the model.. We then measure the matching rate between the output generated from greedy decoding and the outputs from the 100 samples. This matching rate serves as the confidence value for the prompt [^3] This method can approximate the overconfidence calculation in BELIEF of sampling answers from the output distribution. It makes the confidence calculated by BELIEF for encoder-based models comparable to that in BELIEF-ICL.. The calculation of Overconf follows the same setting in §2.4. This method can approximate the BELIEF’s Overconf calculation as BELIEF sampling answers from the output distribution. Note that, due to the high cost of generating 100 samples for each fact, we adopt a more efficient approach. We sample 10K prompts from 10K unique subject-relation pairs and only use these 10K prompts for answer sampling.
@@ 4.1 Dataset construction
We build MyriadLAMA by semi-automatically extending the existing single-prompt probing dataset LAMA-UHN Petroni et al. (2020). MyriadLAMA generates multiple prompts for each fact by providing multiple, equal relational templates for each relation and varying the linguistic expressions of subjects. Additionally, MyriadLAMA offers multiple expressions for each object to cover missed facts that are correctly predicted but in different tokens. For example, for the query “John Lennon was born in [MASK]”, acceptable tokens could include “UK” and “Britain.” [^6] Note that, we follow the setting in LAMA-UHN to include triples where the object is a single token according to the BERT tokenizer. During evaluation, we consider the fact to be present, if the model’s predicted token matches any of the correct tokens, regardless of which correct answer is predicted.
@@ 4.2 Dataset Statistics
We report the statistics of LAMA-UHN and MyriadLAMA, as shown in Table 1. As the result, the number of derived triples is increased from 27,106 in LAMA-UHN to 21,140,500, by combining various semi-automatically generated relational templates and the alias expressions for subject and object entities. As the prompts are generated from derived triples without considering the object expressions, the number of generated prompts are less than the number of derived triples, which is increased from 24,643 to 6,492,800.
Please refer to the appendix for details on dataset construction (§ A) and validity analysis of MyriadLAMA (§ C). Examples of extended templates are provided in § A.3.
@@ 5.1 Experimental setups
LLMs (#params) Pre-training corpora
@@
[…]
BERTlarge (336M) 3.3G
BERTwwm (336M) 3.3G
[…]
Llama2-13B (13B) 2.0T
Llama2-70B (70B) 2.0T
[…]
Llama3-70B (70B) 15T+
[…]
Phi3-small (7B) 4.8T
Phi3-medium (14B) 4.8T
@@
We introduce the encoder- and decoder-based models used in experiments and evaluation settings. We evaluated three encoder-based models, BERTbase, BERTlarge, and BERTwwm, [^9] BERTwwm masks all tokens for a single word at the same time, while BERTbase and BERTlarge masks a single token. and eight decoder-based models, Llama2-7B, -13B, -70B, Llama3-8B, -70B, Phi3-mini, -small, and -medium (Table 2). Refer to §F for more detailed model information.
We conduct a full-scale evaluation on LLMs with up to 8 billion parameters. Due to the computational cost of LLM inference, we use only five manually rewritten templates for the LLMs with more than 8B parameters, including Llama2-70B, Llama3-70B, and Phi3-medium. [^10] The partial evaluation is sufficient to compare performance across different model sizes. To calculate the average and fluctuation of accuracy (§2.2), we set a large sample number ($N=50,000$) to provide stable and accurate accuracy measurement.
@@ 5.2 Do BELIEFs provide additional insights?
BELIEF can offer evaluation from diverse perspectives rather than accuracy. As shown in Table 3 (Above), the evaluation result highlights accuracy fluctuations among the BERT variants. All BERT models show low consistency and tend to be overconfident in their predictions. Figure 1 (left) depicts the relationship between confidence and Acc@1 of the BERT models, indicating low accuracy even for prompts with confident outputs. Whereas BERTwwm performs better over most BELIEF metrics, BERTlarge outperforms BERTwwm on LAMA-UHN. This discrepancy arises from the limited prompts used in LAMA-UHN and the the single-faceted evaluation method. This highlights BELIEF’s effectiveness in achieving a more accurate factual probing comparison between PLMs.
@@ 5.3 Does ICL adhere to instructions?
We then explore the effectiveness of different ICL settings in extracting facts from LLMs. We evaluate the instruction adherence of these settings from two aspects: predicting facts and generating one-word answers, which are crucial since the target objects in MyriadLAMA are primarily one-word entities. We assessed the fact prediction and one-word generation abilities of Llama2-7B, Llama3-8B, and Phi3-small using average Acc@1 and the rate of one-word generation. As shown in Table 4, only under few-shot settings, LLMs demonstrate a remarkable ability to understand instructions for fact prediction and generating one-word answers, indicating the effectiveness of prompting LLMs to predict mask tokens through in-context learning and the risk of evaluating LLMs under zero-shot settings. Our evaluation with a QA-style ICL settings also confirm the same insight (see § D for details). Moreover, using exemplars similar to the target prompt in the few-shot context can achieve improvement overall metrics (Table 3 Below), including the one-word generation rate (Table 4).
ICL settings Fact prediction (Average Acc@1) one-word generation
@@ 5.4 Can BELIEFs mitigate bias?
We explore whether BELIEF can mitigate prompt bias in evaluations. To measure prompt bias quantitatively, we use content-free prompts, where the subject is replaced by meaningless tokens Zhao et al. (2021); Xu et al. (2024), and collect the probabilities of candidate tokens in the output distributions over the mask token. [^12] Specifically, we adopt a similar setting to  Zhao et al. (2021), by ensembling the distribution over prompts with three content-free tokens: “N/A,” an empty string, and “?”. We measure the bias level of the prompt using the certainty of distributions over candidate tokens. Specifically, we define bias level as follows:

### v3

# What Matters in Memorizing and Recalling Facts? Multifaceted Benchmarks for Knowledge Probing in Language Models

@@ Abstract
Language models often struggle with handling factual knowledge, exhibiting factual hallucination issue. This makes it vital to evaluate the models’ ability to recall its parametric knowledge about facts. In this study, we introduce a knowledge probing benchmark, BELIEF(-ICL), to evaluate the knowledge recall ability of both encoder- and decoder-based pre-trained language models (PLMs) from diverse perspectives. BELIEFs utilize a multi-prompt dataset to evaluate PLM’s accuracy, consistency, and reliability in factual knowledge recall. To enable a more reliable evaluation with BELIEFs, we semi-automatically create MyriadLAMA, which has massively diverse prompts. We validate the effectiveness of BELIEFs in comprehensively evaluating PLM’s knowledge recall ability on diverse PLMs, including recent large language models (LLMs). We then investigate key factors in memorizing and recalling facts in PLMs, such as model size, pretraining strategy and corpora, instruction-tuning process and in-context learning settings. Finally, we reveal the limitation of the prompt-based knowledge probing. The MyriadLAMA is publicized. [^1] https://huggingface.co/datasets/iszhaoxin/MyriadLAMA
What Matters in Memorizing and Recalling Facts?
Multifaceted Benchmarks for Knowledge Probing in Language Models
Xin Zhao The University of Tokyo xzhao@tkl.iis.u-tokyo.ac.jp                        Naoki Yoshinaga   Daisuke Obathanks: Currently, he works for ELYZA, Inc. Institute of Industrial Science, The University of Tokyo {ynaga,oba}@iis.u-tokyo.ac.jp
@@ 1 Introduction
One of the strongest motivations for training a language model (LM) using massive text is to increase the ability to handle factual knowledge Kamalloo et al. (2023). However, even if LMs are trained on massive text, they suffer from hallucinations that generate incorrect knowledge-grounded sentences Zhang et al. (2023). Considering that large LMs (LLMs) are being widely applied to real-world tasks, it is vital to evaluate the ability to recall the LLMs’ parametric knowledge and what factors influence on memorizing facts during pre-training.
@@
However, evaluating the LLM’s knowledge recall ability is still challenging. Although LAMA probe Petroni et al. (2019) evaluates the knowledge stored in pre-trained LMs (PLMs), it provides only prediction accuracy. Some studies diversify prompts in the LAMA probe to compute prediction consistency (robustness) Elazar et al. (2021); Jiang et al. (2020), but those datasets have either low quality or low quantity issues (§B). Moreover, since the LAMA probe assumes encoder-based PLMs with the masked LM objective to solve fill-in-the-blank tasks, directly applying it to decoder-based LLMs will underestimate their knowledge recall ability. Although recent studies leveraged QA datasets to probe LLMs’ knowledge Kalo and Fichtel (2022); Mallen et al. (2023); Wiland et al. (2024); Maekawa et al. (2024), they overlook other important aspects than prediction accuracy such as robustness to diverse prompts and the reliability of predictions, which are important for real-world applications.
Figure 1: BELIEFs with MyriadLAMA: the BELIEF benchmarks utilize diverse factual prompts (here, MyriadLAMA) to assess LM’s knowledge recall ability in terms of accuracy, consistency, and reliability.
@@
In this study, we introduce a multifaceted benchmark for knowledge probing, BELIEFs (Figure 1), including BELIEF (§2) and BELIEF-ICL (§3) for encoder- and decoder-based PLMs. BELIEFs utilize diverse prompts for each fact to account for the impact of linguistic expressions when evaluating LLMs’ knowledge recall ability. This allows us to evaluate the robustness and reliability of LLM knowledge by measuring fluctuations in accuracy, consistency, and overconfidence in fact prediction. Since BELIEFs require a multi-prompt probing dataset with diverse prompts for each fact, we build a new probing dataset, MyriadLAMA, to enable a more accurate and comprehensive evaluation (§4). MyriadLAMA expands LAMA-UHN Petroni et al. (2020) by offering different prompts for each fact through a semi-automatic method. Specifically, we obtain a wide variety of lexically, syntactically, and semantically diverse prompts by rewriting relational templates and extending subject expressions.
@@
We applied BELIEFs to various encoder- and decoder-based PLMs, including BERT Devlin et al. (2019) and Llama3 Dubey et al. (2024)5.1). Through extensive evaluations, we verify the utility of BELIEFs in uncovering PLMs’ knowledge recall ability (§5). Moreover, by comparing different PLMs, we gain insights into the factors affecting knowledge recall of PLMs from three aspects: accuracy, reliability and robustness (§6).
The primary findings in this study are as follows:
  • Model size, pretraining strategy, and corpora are crucial factors for memorizing knowledge in LMs during pretraining.
  • Whereas instruction-tuning enhances LLMs’ ability to follow instructions in BELIEF-ICL, it reduces their knowledge recall ability.
  • The inclusion and selection of demonstrations impact knowledge recall, revealing the gap between memorized and recallable facts.
  • Exploring the upper limits of covered knowledge by various methods reveals the limitation of prompt-based knowledge probing (§7).
2 BELIEF Benchmark
@@
We first present the multifaceted factual probing benchmark, BELIEF for encoder-based PLMs. Using a multi-prompt probing dataset, BELIEF evaluates the knowledge recall ability of PLMs from accuracy, robustness and reliability (§2.2-2.4). Here, robustness measures PLMs’ ability to maintain consistent accuracy and predictions when given different prompts in evaluation. Reliability reflects the extent to which we can trust the PLMs’ predictions.
2.1 Preliminaries
@@
To evaluate the facts in PLMs, BELIEF aggregates results from multiple prompts for each fact to mitigate biases from specific linguistic expressions. This requires varied expressions for each fact, namely multi-prompt factual probing dataset.
@@ 2.2 Accuracy and its fluctuations
Fluctuation of accuracy: For $V_{\textrm{Acc}@1}$, we evaluate accuracy fluctuations using the range and standard deviation (SD). The range is determined by the difference between the maximum and minimum accuracy values in $V_{\textrm{Acc}@1}$.
@@ 2.4 Reliability
Ovconf $=\sum_{i=1}^{M}\frac{|P^{(i)}|}{M}(\overline{o_{max}}^{(i)}-\overline{\textrm{Acc}@1}^{(i)})$ (2)
The closer the Ovconf is to zero, the more aligned the model’s confidence is with its accuracy, indicating reliable confidence. A negative Ovconf value means the model is underconfident.
@@ 3.1 In-context learning for fact probe
2) Context settings: We propose four types of contexts to assess the impact of examplar selection on factual knowledge probing, following the QA format outlined in InstructGPT Ouyang et al. (2022). zero-shot uses only instructions; X-random samples X facts from all relations as the few-shot demonstrations; X-relation samples X facts from the same relation but with random templates; X-template samples X facts from the same relations and the same template.
@@ 3.2 Evaluation methods
To measure BELIEF metrics for LLMs, we compare two strings: the generated text and the correct object expression for Acc@1, and two generated texts for Consist and Ovconf. Here, we first normalize strings by tokenizing and lemmatizing them. For example, “a guitar” and “guitars” are normalized to “a, guitar” and “guitar.” If a normalized string list is included in the other (partial matching), they are considered matched.
@@
3) Reliability: To calculate overconfidence, we need the model’s confidence (probability) in its output. However, we cannot obtain this directly from the probability over generated tokens, as LLMs can produce diverse outputs that represent the same answer. To address this, we propose an approximate measurement. For each prompt, we generate 100 samples using multinomial sampling [^2] Multinomial sampling selects a next token according to the probability, over the entire vocabulary given by the model.. We then measure the matching rate between the output generated from greedy decoding and the outputs from the 100 samples. This matching rate serves as the confidence value for the prompt [^3] This method can approximate the overconfidence calculation in BELIEF of sampling answers from the output distribution. It makes the confidence calculated by BELIEF for encoder-based models comparable to that in BELIEF-ICL.. The calculation of Ovconf follows the same setting in §2.4. This method can approximate the BELIEF’s Ovconf calculation as BELIEF sampling answers from the output distribution. Note that, due to the high cost of generating 100 samples for each fact, we adopt a more efficient approach. We sample 10K prompts from 10K unique subject-relation pairs and only use these 10K prompts for answer sampling.
@@ 4.1 Dataset construction
We build MyriadLAMA by semi-automatically extending the existing single-prompt probing dataset LAMA-UHN Petroni et al. (2020). MyriadLAMA generates multiple prompts for each fact by providing multiple, equal relational templates for each relation and varying the linguistic expressions of subjects. Additionally, MyriadLAMA offers multiple expressions for each object to cover missed facts that are correctly predicted but in different tokens. For example, for the query “John Lennon was born in [MASK]”, acceptable tokens could include “UK” and “Britain.” [^6] We follow the setting of LAMA-UHN triples where the object is a single token according to the BERT tokenizer. During evaluation, we consider the fact to be present, if the model’s predicted token matches any of the correct tokens, regardless of which correct answer is predicted.
@@ 4.2 Dataset Statistics
Table 1 lists the statistics of MyriadLAMA. The number of derived triples is increased from 27,106 in LAMA-UHN to 21,140,500, by combining various semi-automatically generated relational templates and the alias expressions for subject and object entities. As the prompts are generated from derived triples without considering the object expressions, the number of generated prompts are less than the number of derived triples, which is increased from 24,643 to 6,492,800. Refer to the appendices for details on dataset construction (§A) and validity analysis of MyriadLAMA (§C). Examples of extended templates are provided in §A.3.
@@ 5.1 Experimental setups
We use BELIEFs to evaluate the knowledge recall abilities of both encoder- and decoder-based PLMs. The target encoder-based PLMs include BERTbase, BERTlarge, and BERTwwm. [^9] BERTwwm masks all tokens for a single word at the same time, while BERTbase and BERTlarge masks a single token. The target decoder-based LLMs include Llama2 (7B, 13B, and 70B) and Llama3 (8B and 70B), without and with instruction tuning (except for Llama3-70B), along with Phi3 (mini, small, and medium). Their brief pretraining information are listed in Table 2. Refer to §F for more details.
PLMs (#params) Pre-training corpora
@@
BERTbase (110M) $\left.\begin{array}[]{@{}l@{}}\text{3.3B+}\\ \text{3.3B+}\\ \text{3.3B+}\end{array}\right\}$ English Wikipedia & BookCorpus
BERTlarge (336M)
BERTwwm (336M)
Llama2-7B(-IT) (7B) $\left.\begin{array}[]{@{}l@{}}\text{2.0T}\\ \text{2.0T}\\ \text{2.0T}\end{array}\right\}$ A collection of publicly available online data.
Llama2-13B(-IT) (13B)
Llama2-70B(-IT) (70B)
Llama3-8B(-IT) (8B) $\left.\begin{array}[]{@{}l@{}}\text{15T+}\\ \text{15T+}\end{array}\right\}$ A collection of publicly available online data.
Llama3-70B (70B)
Phi3-mini (3.8B) $\left.\begin{array}[]{@{}l@{}}\text{4.9T}\\ \text{4.9T}\\ \text{4.9T}\end{array}\right\}$ High-quality educational data/code/chat & synthetic textbook-like data
Phi3-small (7B)
Phi3-medium (14B)
@@
We conduct a full-scale evaluation on LLMs with up to 8 billion parameters. To save cost of LLM inference, we use five manually rewritten templates only for the LLMs with more than 8B parameters, including Llama2-70B and its IT variant Llama2-70B-IT, Llama3-70B, and Phi3-medium. [^10] The partial evaluation is sufficient to compare performance across different model sizes. To calculate the average and fluctuation of accuracy (§2.2), we set a large sample number ($N=50,000$) to provide stable, accurate result.
@@ 5.2 Do BELIEFs provide additional insights?
BELIEFs offer evaluation from diverse perspectives rather than accuracy. As shown in Table 3 (Above), the evaluation result highlights accuracy fluctuations among the BERT variants. All BERT models show low consistency and tend to be overconfident in their predictions. Figure 2 (left) depicts the relationship between confidence and Acc@1 of the BERT models, indicating low accuracy even for prompts with confident outputs. Whereas BERTwwm performs better over most BELIEF metrics, BERTlarge outperforms BERTwwm on LAMA-UHN. This discrepancy arises from the limited prompts used in LAMA-UHN and the the single-faceted evaluation method. This highlights BELIEF’s effectiveness in achieving a more accurate factual probing comparison between PLMs.
@@ 5.3 Does ICL adhere to instructions?
We then explore the effectiveness of different ICL settings in extracting facts from LLMs. We evaluate the instruction adherence of these settings from two aspects: predicting facts and generating one-word answers, reflecting that the target objects in MyriadLAMA are primarily one-word entities.
Table 4 shows Acc@1 and one-word generation ratio of two pretrained LLMs (Llama2-7B and Llama3-8B) and one instruction-tuned LLM (Phi3-small). We found that under few-shot settings, even the pretrained LLMs exhibit a remarkable ability to follow instructions, indicating the effectiveness of prompting LLMs to predict mask tokens through in-context learning. Our evaluation with QA-style ICL settings also confirms this (see §D for details). Moreover, exemplars similar to the target prompt (4-template) in the context boosted improvement overall metrics (Table 3 Below, Table 4).
ICL settings Fact prediction (Acc@1) 1-word ratio
@@ 5.4 Can BELIEFs mitigate bias?
We explore whether BELIEFs can mitigate prompt bias in evaluations. To measure prompt bias quantitatively, we use content-free prompts, where the subject is replaced by meaningless tokens Zhao et al. (2021); Xu et al. (2024), and collect the probabilities of candidate tokens in the output distributions over the mask token. [^12] Specifically, we adopt a similar setting to  Zhao et al. (2021), by ensembling the distribution over prompts with three content-free tokens: “N/A,” an empty string, and “?”. We measure the bias level of the prompt using the certainty of distributions over candidate tokens. Specifically, we define bias level as follows:
bias-level $=1-\frac{\mathcal{H}}{\mathcal{H}_{max}}$ (3)
We measure bias in both single- and multi-prompt evaluations. In single-prompt evaluation, we represent bias as the average level across all relational templates. For measuring bias-level in multi-prompt evaluation, we first average output distributions by different templates for each relation, then use the bias-level of the averaged distribution to quantify it. Taking P31:instance-of [^13] https://www.wikidata.org/wiki/Property:P31 as an example, the average probability of “science” over all templates is 8.30%, but it rises to 52.79% for template: “[Y] contains [X] as one of its elements.”
6 Differentiating PLMs in Fact Probing
This section compares the PLMs’ knowledge recall abilities in terms of accuracy, reliability, and robustness and then explores factors affecting them.
6.1 Factors affecting the recall accuracy
1) Pre-training strategy. Table 3 confirms that BERTwwm outperforms BERTlarge in terms of all metrics, while BERTwwm differs from BERTlarge only in the masking strategy during pre-training. The superiority of BERTwwm likely stems from its challenging pre-training paradigm, which requires recalling whole words without sub-token information, enhancing word-level contextual understanding. This underscores the importance of pre-training strategy in knowledge acquisition.